One number is 11 more than twice another number. If the sum of the numbers is twice their difference, find the numbers
Answer by stanbon(75887)

One number is 11 more than twice another number. If the sum of the numbers is twice their difference, find the numbers
-----
Equations:
x = 2y + 11
x+y = 2(x-y)
-------------
Simplify the 2nd equation:
x+y = 2x-2y
x = 3y
---
Substitute that into the 1st equation:
3y = 2y+11
y = 11
----
Solve for "x":
x = 3y
x = 3*11
x = 33
============
Cheers,
Stan H.
=============